# Higher property T and below-rank phenomena of lattices

### v1

@@ 3.4.2. Degree 2: Stability
Cohomological vanishing in degree 2 and Banach coefficients has applications in the theory of stability, see [CGLT]. We will not discuss it here in length, as this topic was covered in Thom’s excellent manuscript [Thom]. Here we will only make some comments on recent developments. [Thom, Question 3.10] was answered positively in [badsau], see Theorem 1. [Thom, Question 3.11] was answered positively for $1<p<\infty$ in Lubotzky-Oppenheim [Lub-Opp], using vanishing of second cohomology in certain $L^{p}$-representations to show Schatten-$p$-norm stability and non-approximability. For every $1<p_{1}<p_{2}<\infty$ they constructed groups with vanishing of second cohomology with coefficients in these $L^{p}$ representations, for every $p\in[p_{1},p_{2}]$. They used that to construct a finitely presented group which is not $p$-norm approximated for any $1<p<\infty$. They notably left open the cases $p=1$ and $p=\infty$, both with respect to cohomological vanishing and non-approximability. The non-approximability question in $p=1$ was very recently solved by Bachner-Dogon-Lubotzky in [BDLub], in a way bypassing the cohomological vanishing question. Corollary 8 settles completely the question of cohomological vanishing also in the case $p=1$. The case $p=\infty$ remains open. We note, however, that the discouraging reasoning given by the end of the introduction of [Lub-Opp], describing why their methods must be insufficient for this case, seems somewhat less discouraging in view of Theorem 45 and Conjecture 47 which regard vanishing of cohomology with von Neumann algebra coefficients.

### v2

@@ 3.4.2. Degree 2: Stability
An excellent background reference for stability is Thom’s article [96]. Here we will only make some comments on recent developments.
Cohomological vanishing in degree 2 has applications in the theory of stability. This was an important insight of the work of de Chiffre-Glebsky-Lubotzky-Thom [36]. Based on this and Garland’s work they constructed central extensions of certain higher rank $p$-adic lattices that are not Frobenius-approximable. Based on [36] and on [15], Bader-Lubotzky-Sauer-Weinberger showed that the non-trivial central extensions of symplectic lattices are not Frobenius-approximable [12].
[96, Question 3.10] was answered positively in [15], see Theorem 1. [96, Question 3.11] was answered positively for $1<p<\infty$ in Lubotzky-Oppenheim [74], using vanishing of second cohomology in certain $L^{p}$-representations to show Schatten-$p$-norm stability and non-approximability. For every $1<p_{1}<p_{2}<\infty$ they constructed groups with vanishing of second cohomology with coefficients in these $L^{p}$ representations, for every $p\in[p_{1},p_{2}]$. They used that to construct a finitely presented group which is not $p$-norm approximated for any $1<p<\infty$. They notably left open the cases $p=1$ and $p=\infty$, both with respect to cohomological vanishing and non-approximability. The non-approximability question in $p=1$ was very recently solved by Bachner-Dogon-Lubotzky in [6], in a way bypassing the cohomological vanishing question. Corollary 12 settles completely the question of cohomological vanishing also in the case $p=1$. The case $p=\infty$ remains open. We note, however, that the discouraging reasoning given by the end of the introduction of [74], describing why their methods must be insufficient for this case, seems somewhat less discouraging in view of Theorem 51 and Conjecture 53 which regard vanishing of cohomology with von Neumann algebra coefficients.
3.4.3. Degree 2: Applications to measurable diversity
The fact that there are hyperbolic groups, like uniform lattices in the isometry group of the Cayley hyperbolic plane, with highter property T (see Theorem 5) combined with the theory of group-theoretic Dehn fillings [82, 35], especially work of Petrosyan–Sun [93, 86, 87], is an interesting source of tension between rigidity and flexibility phenomena. In [45] Fornier-Facio and the second author use the cohomological excision theorem for Cohen-Lyndon triples and property $(T_{2})$ to obtain a good control on the second cohomology of hyperbolic quotients obtained by group-theoretic Dehn fillings. See also [46] for a similar construction that shows that poperty $(T_{2})$ does not pass to quotients in general. The rigid control on the second cohomology together with the flexibility of group-theoretic Dehn fillings allows to construct an uncountable family consisting of colimits of quotients of a fixed hyperbolic group with property $(T_{2})$ such that the range of the second $\ell^{2}$-Betti number is uncountable. More precisely, the following can be shown if one takes a uniform lattice in the isometry group of the Cayley hyperbolic plane as a starting point.
Theorem 115 ([45]*Theorem A).
There exists an uncountable family of simple Kazhdan groups of rational cohomological dimension $16$ with pairwise distinct positive second $\ell^{2}$-Betti number. Furthermore, these groups are pairwise non-measure equivalent.
It is an easy consequence of Gaboriau’s construction [51]*Proposition VI.16 that there exist uncountably many pairwise non-measure equivalent groups [63]*Proposition 5.1. The previous theorem is the first construction of a measurably diverse family with bounded rational cohomological dimension. Two constructions of measurably diverse families of Kazhdan groups have appeared last year and precede the above theorem. One is due to Fournier-Facio and Sun [44], where the groups are distinguished by their entire sequence of $\ell^{2}$-Betti numbers, in a spirit similar to the work of López Neumann [71], who constructs an infinite family of finitely presented simple groups. Very recently, Ioana–Tucker-Drob [63] produced further examples of groups with vanishing $\ell^{2}$-Betti numbers, arising from the theory of wreath-like products [32]. In particular, these groups have infinite rational cohomological dimension.